OpenStudy (anonymous):

Solve the absolute value equation. Graph the solution. │3x + 1│ = 1 x = 1 or x = x = or x = 0 x = 1 or x = –2 x = 1 or x = 0

OpenStudy (anonymous):

@Butterfly16

OpenStudy (anonymous):

Set up two equations, because the absolute value of both -1 and 1 is 1 so 3x+1=1 and 3x+1=-1

OpenStudy (anonymous):

│3x + 1│ = 1 (3x+1)=1 and -(3x+1)=1 Can you figure it out? :o)

OpenStudy (klimenkov):